OpenStudy (anonymous):

Find the three arithmetic means between -2 and 10. Answer 2, 4, 7 1, 5, 7 1, 4, 7 2, 5, 8

OpenStudy (amistre64):

divide the range into 4 equal parts ... the rest is addition

OpenStudy (amistre64):

arithmetic, not geometric ... a common difference

OpenStudy (anonymous):

how do i set that up?

OpenStudy (amistre64):



OpenStudy (amistre64):

-2 + 4d = 10, what is d?

OpenStudy (anonymous):

so 3?

OpenStudy (amistre64):

yep, so we simply add 3 to each term

OpenStudy (amistre64):

-2+3 -2+3+3 10-3 ... or -2+3+3+3 should be the 3 terms wanted.

OpenStudy (anonymous):

so the answer would be 1,4,7 correct?

OpenStudy (amistre64):

correct :)

OpenStudy (anonymous):

thank you so much@!!!:D

OpenStudy (anonymous):

Find S22 for -6 + (-4) + (-2) + 0 … how do i attack this?

OpenStudy (amistre64):

theres a formula we could develop, but its usually just simpler to apply the concept instead of memorizing the formula good luck :)

OpenStudy (amistre64):

S22 is the sum of the first 22 terms of a sequence that generates the values

OpenStudy (amistre64):

looks arithmatic to me: a0 = -6+2(0) a1 = -6+2(1) a2 = -6+2(2) ... an = -6+2n